• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name weinstein

    Log entry time 12:01:24 on June30,2000

    Entry number 45302

    keyword=Shift summary


    Shift Summary 6/30 8-16
    Did optics scan with sieve slit
    Did not do pointing studies

    8:00 waiting for beam to return well centered
    8:10 start run 2678 H(e,e) sieve at delta = -4%
    8:13 Q3 trip - overtemperature fault - Ed Folts coming in
    8:40 resetting Q3, cycling to 1600 A
    9:00 Q3 back at 1515 A, waiting for Hall to be secured
    9:20 beam back, start run 2679
    9:45 go to -2%
    9:59 sstart run 2680
    10:20 change to 0% in HRSe, put in hadron 6msr coll
    10:32 start run 2681, delta = 0% and q pointing, ps set to 1/alot/alot/alot/1
    OOPS! discovered that the ps1 had been 23 from 8 to 10:32
    10:45 start run 2682, H(e,e'p) elastic with 6msr in each arm
    10:48 increase ps1 from 3 to 10 and start run 2683
    10:52 increase ps1,3,5 to 20/6/2 and start run 2684
    11:05 change to C tgt, C(e,e'p) quasielastic run 2685
    11:18 change to H(e,e) sieve slit in HRSe, LH2, delta = +2%
    11:29 start run 2686
    11:34 end run, change field to +4%, p_e = 3.5136
    11:44 start run 2687
    11:48 goto 6msr in HRSe, 12C, 5uA, no raster
    11:57 start run 2688
    11:57:23 RF trip, goodbye beam